40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 14 09:28:46 2021, Yehonathan, Update, PSL, PMC unlocked 
    Reply  Thu Oct 14 11:25:49 2021, Yehonathan, Update, PSL, PMC unlocked BS_OSEM_Sensor_PSD.pdfBS_OSEM_Sensor_PSD_AfterReconnectingCables.pdf
Message ID: 16400     Entry time: Thu Oct 14 09:28:46 2021     Reply to this: 16401
Author: Yehonathan 
Type: Update 
Category: PSL 
Subject: PMC unlocked 

PMC has been unlocked since ~ 2:30 AM. Seems like the PZT got saturated. I moved the DC output adjuster and the PMC locked immidiatly although with a low transmission of 0.62V (>0.7V is the usual case) and high REFL.

IMC locked immidiately but IFO seems to be completely misaligned. The beams on the AS monitor are moving quite alot syncronously. BS watchdog tripped. I enabled the coil outputs. Waiting for the RMS motion to relax...

Its not relaxing. RMS motion is still high. I disabled the coils again and reenabled them. This seems to have worked. Arms were locked quite easily but the ETMs oplevs were way off and the ASS couldn't get the TRX and TRY more than 0.7. I align the ETMs to center the oplev. I realign everything else and lock the arms. Maximium TR is still < 0.8.

 

 

ELOG V3.1.3-